# High Gain Fusion Target Design using Generative Artificial Intelligence

### v4

@@ Abstract
By returning to the topological basics of fusion target design, Generative Artificial Intelligence (genAI) is used to specify how to initially configure and drive the optimally entangled topological state, and stabilize that topological state from disruption. This can be applied to all methods; including tokamaks, laser-driven schemes, and pulsed-power driven schemes. The result is practical, room temperature targets that can yield up to 10 GJ of energy, driven by as little as 3 MJ of absorbed energy. The genAI is based on the concept of Ubuntu that replaces the Deep Convolutional Neural Network approximation of a functional, with the formula for the generating functional of a canonical transformation from the domain of the canonical field momentums and fields, to the domain of the canonical momentums and coordinates, that is the Reduced Order Model. This formula is a logical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix, given observation of the fields. This can be viewed as topological characterization and control of collective, that is complex, systems.
@@ I Return to Topological Basics
The task of fusion target and drive design is to optimize the entanglement of the topological state and stabilize the topological state from disruption. Up to now, there has not been a logical mathematical process via which this can be done. Generative Artificial Intelligence, based on the concept of Ubuntu, provides the logical mathematical process. The African philosophy of Ubuntu is based on “interconnectedness”, that leads to a conservative system with canonical structure, that is a symplectic geometry. Given this logical mathematical process, that is Ubuntu genAI, the collective system can be controlled – the topology of the plasma can be optimized and stabilized.
@@
Instead of a cylindrical (2D) implosion for ignition and cylindrical (2D) burn, Ubuntu Fusion Target Technology gives a more efficient spherical (3D) implosion and spherical to cylindrical (2.5D) burn. Instead of a constraint on linear growth, Ubuntu Fusion Target Technology provides the freedom of nonlinear self-organization that leads to a 2D implosion for fuel assembly and a 3D implosion for ignition. This leads to a fundamentally different and much more optimistic scaling than that of Betti [12], Schmit, and Ruiz [13]. Their scaling and design philosophy is based on the suppresion of the linear growth of plasma instability; where Ubuntu Fusion Target Technology is based on the optimization and stabilization of the nonlinear topological plasma states – being a good parent and encouraging the plasma to mature and become its own person.
@@
Figure 3: The Ubuntu laser-driven fusion target design: (a) a Be cylindrical shell, coated on the inside with solid fuel, and filled with DT gas; shown is the Z-pinch laser drive, a DALL-E rendered liner plasma tornado, an experimental back-lit image from the University of Michigan of the liner plasma to the right of the plasma tornado, and a self emission image of the DD MagLIF plasma stagnation in orange, and (b) the twisted pair of the Z-pinch stagnation topology, with the return current.
@@ II The Ubuntu Fusion Target
For brevity, we will describe the laser targets and drives. The Ubuntu Fusion Target is a simple, easy to fabricate, room temperature target. This is in contrast to the very impractical and complicated MagLIF DT ice burner, and current ICF target designs. As shown in Fig. 3, the target consists of a cylinder of beryllium, coated on the inside with solid fuel, and filled with DT gas. The target is driven by four lasers whose illumination patterns on the cylinder are shaped with phase plates, as shown in Fig. 4, either to form a Z-pinch or an RFP. The pattern of the laser drive effectively makes the target plasma into a substantial MIFED coil. The implosion is not driven by radiation pressure nor ablation pressure, but by magnetic pressure which increases as the target is imploded. Unlike pulsed-power, which must physically touch the plasma, the lasers stand off from the target and only require four small lines of sight in a plane perpendicular to the axial direction. Ubuntu genAI will be used to optimize the target and drive design, including the axial structure. It will also be used to ponderomotively stabilize the metastable stagnation.
@@ III Ubuntu Generative Artificial Intelligence
where $z\equiv p+\text{i}\,x$, $\phi$ is a father wavelet of scale $1/p$ and position $x$ or pooling operator, $G(z)=R(z)=\text{i}\ln(R_{0}(z))$ is the activation function, $R_{0}(z)=h^{-1}(2z/\pi)/\text{i}$, $h(z)=(z+1/z)/2$, $\psi\star$ is a convolutional layer defined by the mother wavelet of scale $1/p_{k}$, and $p=\sum{p_{k}}$. This functional is a non-stationary, that is local, Fourier transformation. This functional has a fast, that is N-log-N scaling, forward and inverse.
@@
The evidence that led to this new theory was: the work of Mallat on the Scattering Transformation (MST) [20], the Wavelet Phase Harmonics (WPH) [21], and the Wavelet Conditional Renormalization Group (WCRG) [22]; the form of Deep Q-networks (that is, the approximate Q-function, as a solution to the Bellman equation) [17]; the form of GPTs (that is, the approximate score function or approximate log-likelihood, as action or entropy) [18]; and the work of Glinsky and Maupin [11].
@@ IV A Mathematically Logical Process of Renormalization
Dirac was very dissatisfied in his later life with the state of field theory [23]. Like Einstein, he felt that field theory should have a reality, and be a question, not of probability, but of geometry (manifold curvature and geodesic motion) or topology or group symmetry. He also felt that “renormalization is not a logical mathematical process”. The HST is a logical mathematical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix. A Principle Components Analysis (PCA) of the HST gives: the renormalization spectrums, that is the solutions to the Renormalization Group Equations (RGEs); Heisenberg’s Scattering Matrix (S-Matrix) [24]; the $m$-Body Greens functions, which is obvious given the form of $G(z)=\text{i}\ln(R_{0}(z))$ in the formula for the HST, which is the well-known Greens function of a complex variable [25]; the $m$-body scattering cross sections; the Wigner-Weyl transformation [26, 27]; the Mayer Cluster Expansion [28]; or the functional Taylor expansion of the action functional, $\delta^{m}S/\delta f^{m}$. This enables the unification of the four forces as a question of Lie group symmetries, that is a Grand Unified Theory (GUT) of fields, or geodesic motion on the dynamical manifold. This can be seen by Murray Gell-Mann’s Standard Model [29] (where U(1) is the Lie group symmetry of the ElectroMagnetic force, SU(2) is the Lie group symmetry of the weak force, and SU(3) is the Lie group symmetry of the strong force); and Dirac writing down the Hamiltonian, that is the infinitesimal Lie group generator, for Einstein’s General Theory of Relativity [30], in 1958 [31].
@@ V Control of Collective Systems
Collectives are fields or swarms of elementary particles (making elementary fields), charged particles (making plasmas), molecules (making fluids), celestial bodies (making cosmoses), economic entities (making economies), persons (making societies) and so on. This new theory gives a way of controlling (that is, optimizing and stabilizing) collective systems, whose state is given, not by variables, but as a function, $f(x)$. This also can be viewed as topological characterization and control. The two-stage canonical auto-encoder is finding the analytic function; whose Riemann surface (that is, manifold), has the topology of the collective system; and whose non-stationary singularity spectrums are the Riemann moduli or topological indices (that is, the integrals of the Chern-Simons 3-form) of the collective system.
@@
How does Ubuntu genAI control the collective system? First, an application of the renormalization spectrums ponderomotively stabilizes the collective system. This is a critical issue for tokamaks: the tokamak H-mode needs to be stabilized from disruption. Second, the new theory gives very fast surrogate models, with high fidelity (that is topological and Noether invariant conservation). This enables Bayesian optimal design and Bayesian data analysis. For instance, Glinsky and Maupin [11] were able to train an MHD surrogate, using a preliminary version of this new theory, in 20 core*sec using 500 2D simulations. The high fidelity surrogate could simulate the system in 1 core*sec, where traditional finite element methods would take 240 core*hrs. For a 3D simulation, traditional methods would take 100,000 core*hrs. The Singular Value Decomposition of the cross-variance matrix showed a one-to-one mapping between the PCAs, that is the renormalization spectrums, and the initial conditions – a remarkable result that gives physical meaning to the renormalization spectrums.
@@ Acknowledgements.
Thanks is given to CSIRO for supporting the early parts of this research through their Science Leaders Program, and the Institute des Hautes Etudes Scientifique (IHES) for hosting a stay where the initial details of this theory were formalized. Thanks is also given to both Stéphane Mallat and Joan Bruna for many useful discussions, and communication of many of their mathematical results before they were presented or published. The mentorship of both Ted Frankel and Michael Freedman in topology and the geometry of physics is the foundation on which this work was constructed. Finally, thanks is given to the University of Western Australia, John Hedditch, and Ian MacArthur for their help in understanding many of the finer points of Quantum Field Theory.

### v5

@@ Abstract
By returning to the topological basics of fusion target design, Generative Artificial Intelligence (genAI) is used to specify how to initially configure and drive the optimally entangled topological state, and stabilize that topological state from disruption. This can be applied to all methods; including tokamaks, laser-driven schemes, and pulsed-power driven schemes. This target design philosophy has the potential to yield practical, room temperature targets that could yield up to 10 GJ of energy, driven by as little as 3 MJ of absorbed energy. The genAI is based on the concept of Ubuntu that replaces the Deep Convolutional Neural Network approximation of a functional, with the formula for the generating functional of a canonical transformation from the domain of the canonical field momentums and fields, to the domain of the canonical momentums and coordinates, that is the Reduced Order Model. This formula is a logical process of renormalization, that has the potential to enable Heisenberg’s canonical approach to field theory, via calculation of the S-matrix, given observation of the fields. This can be viewed as topological characterization and control of collective, that is complex, systems.
@@ I Return to Topological Basics
The task of fusion target and drive design is to optimize the entanglement of the topological state and stabilize the topological state from disruption. Up to now, there has not been a logical mathematical process via which this can be done. There has only been the empirical methods of Wilson [5] and tHooft [6]. Generative Artificial Intelligence, based on the concept of Ubuntu, provides the logical mathematical process. The African philosophy of Ubuntu is based on “interconnectedness”, that leads to a conservative system with canonical structure, that is a symplectic geometry. Given this logical mathematical process, that is Ubuntu genAI, the collective system can be controlled – the topology of the plasma can be optimized and stabilized.
@@
Instead of a cylindrical (2D) implosion for ignition and cylindrical (2D) burn, Ubuntu Fusion Target Technology gives a more efficient spherical (3D) implosion and spherical to cylindrical (2.5D) burn. Instead of a constraint on linear growth, Ubuntu Fusion Target Technology provides the freedom of nonlinear self-organization that leads to a 2D implosion for fuel assembly and a 3D implosion for ignition. This leads to a fundamentally different and much more optimistic scaling than that of Betti [14], Schmit and Ruiz [15]. Their scaling and design philosophy is based on the suppresion of the linear growth of plasma instability; where Ubuntu Fusion Target Technology is based on the optimization and stabilization of the nonlinear topological plasma states – being a good parent and encouraging the plasma to mature and become its own person.
@@
Figure 3: The Ubuntu laser-driven fusion target design: (a) a Be cylindrical shell, coated on the inside with solid fuel, and filled with DT gas; shown is the Z-pinch laser drive, a DALL-E rendered liner plasma tornado, an experimental back-lit image from the University of Michigan [8] of the liner plasma to the right of the plasma tornado, and a self emission image of the DD MagLIF plasma stagnation in orange, and (b) the twisted pair of the Z-pinch stagnation topology, with the return current.
@@ II The Ubuntu Fusion Target
For brevity, we will describe the laser targets and drives. The Ubuntu Fusion Target is a simple, easy to fabricate, room temperature target. This is in contrast to the impractical and complicated MagLIF DT ice burner, and current ICF target designs. As shown in Fig. 3, the target consists of a cylinder of beryllium, coated on the inside with solid fuel, and filled with DT gas. The target is driven by four lasers whose illumination patterns on the cylinder are shaped with phase plates, as shown in Fig. 4, either to form a Z-pinch or an RFP. The pattern of the laser drive effectively makes the target plasma into a substantial MIFED coil. The implosion is not driven by radiation pressure nor ablation pressure, but by magnetic pressure which increases as the target is imploded. Unlike pulsed-power, which must physically touch the plasma, the lasers stand off from the target and only require four small lines of sight in a plane perpendicular to the axial direction. Ubuntu genAI will be used to optimize the target and drive design, including the axial structure. It will also be used to ponderomotively stabilize the metastable stagnation.
@@ III Ubuntu Generative Artificial Intelligence
where $z\equiv p+\text{i}\,x$, $\phi$ is a father wavelet of scale $1/p$ and position $x$ or pooling operator, $G(z)=R(z)=\text{i}\ln(R_{0}(z))$ is the activation function, $R_{0}(z)=h^{-1}(2z/\pi)/\text{i}$, $h(z)=(z+1/z)/2$, $\psi\star$ is a convolutional layer defined by the mother wavelet of scale $1/p_{k}$, and $p=\sum{p_{k}}$. This functional is a non-stationary, that is local, spectral transformation. This functional has a fast, that is N-log-N scaling, forward and inverse.
@@
The evidence that led to this theory was: the work of Mallat on the Scattering Transformation (MST) [22], the Wavelet Phase Harmonics (WPH) [23], and the Wavelet Conditional Renormalization Group (WCRG) [24]; the form of Deep Q-networks (that is, the approximate Q-function, as a solution to the Bellman equation) [19]; the form of GPTs (that is, the approximate score function or approximate log-likelihood, as action or entropy) [20]; and the work of Glinsky and Maupin [13].
@@ IV A Mathematically Logical Process of Renormalization
Dirac was very dissatisfied in his later life with the state of field theory [25]. Like Einstein, he felt that field theory should have a reality, and be a question, not of probability, but of geometry (manifold curvature and geodesic motion) or topology or group symmetry. He also felt that “renormalization is not a logical mathematical process”. The HST is a logical mathematical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix. A Principle Components Analysis (PCA) of the HST gives: the renormalization spectrums, that is the solutions to the Renormalization Group Equations (RGEs); Heisenberg’s Scattering Matrix (S-Matrix) [26]; the $m$-Body Greens functions, which is obvious given the form of $G(z)=\text{i}\ln(R_{0}(z))$ in the formula for the HST, which is the well-known Greens function of a complex variable [27]; the $m$-body scattering cross sections; the Wigner-Weyl transformation [28, 29]; the Mayer Cluster Expansion [30]; or the functional Taylor expansion of the action functional, $\delta^{m}S/\delta f^{m}$. This has the potential to unify the four forces as a question of Lie group symmetries, or geodesic motion on the dynamical manifold. This can be seen by Murray Gell-Mann’s Standard Model [31] (where U(1) is the Lie group symmetry of the ElectroMagnetic force, SU(2) is the Lie group symmetry of the weak force, and SU(3) is the Lie group symmetry of the strong force); and Dirac writing down the Hamiltonian, that is the infinitesimal Lie group generator, for Einstein’s General Theory of Relativity [32], in 1958 [33].
@@ V Control of Collective Systems
Collectives are fields or swarms of elementary particles (making elementary fields), charged particles (making plasmas), molecules (making fluids), celestial bodies (making cosmoses), economic entities (making economies), persons (making societies) and so on. This theory gives a way of controlling (that is, optimizing and stabilizing) collective systems, whose state is given, not by variables, but as a function, $f(x)$. This also can be viewed as topological characterization and control. The two-stage canonical auto-encoder is finding the analytic function; whose Riemann surface (that is, manifold), has the topology of the collective system; and whose non-stationary singularity spectrums are the Riemann moduli or topological indices (that is, the integrals of the Chern-Simons 3-form) of the collective system.
@@
How does Ubuntu genAI control the collective system? First, an application of the renormalization spectrums ponderomotively stabilizes the collective system. This is a critical issue for tokamaks: the tokamak H-mode needs to be stabilized from disruption.
Second, the theory gives very fast surrogate models, with high fidelity (that is topological and Noether invariant conservation). This enables Bayesian optimal design and Bayesian data analysis. For instance, Glinsky and Maupin [13] were able to train an MHD surrogate, using a preliminary version of this theory, in 20 core*sec using 500 2D simulations. The high fidelity surrogate could simulate the system in 1 core*sec, where traditional finite element methods would take 240 core*hrs. For a 3D simulation, traditional methods would take 100,000 core*hrs. The Singular Value Decomposition of the cross-variance matrix showed a one-to-one mapping between the PCAs, that is the renormalization spectrums, and the initial conditions – a remarkable result that gives physical meaning to the renormalization spectrums.
Better yet, the two stage canonical autoencoder can be configured to generate an ensemble of the collective system in dynamic equilibrium [21]. This replaces the Bayesian optimal design and the Bayesian data analysis, which are computationally slow, and assume a statistical equilibrium.
@@ Acknowledgements.
Thanks is given to CSIRO for supporting the early parts of this research through their Science Leaders Program, and the Institute des Hautes Etudes Scientifique (IHES) for hosting a stay where the initial details of this theory were formalized. Thanks is also given to both Stéphane Mallat and Joan Bruna for many useful discussions, and communication of many of their mathematical results before they were presented or published. The mentorship of both Ted Frankel and Michael Freedman in topology and the geometry of physics is the foundation on which this work was constructed. Finally, thanks is given to the University of Western Australia, John Hedditch, and Ian MacArthur for their help in understanding many of the finer points of Quantum Field Theory. The subject matter of this Letter is the basis of the US Patent Application US20250116976-A1, entitled ”Systems and Methods for Controlling Complex Systems”, and the Provisional US Patent Application 63/760917, entitled ”Inertial Fusion Energy Generation Systems and Methods”.